# Epidemiology-informed Graph Neural Network for Heterogeneity-aware Epidemic Forecasting

### v1

@@ VI Conclusion
In this study, we propose a novel epidemic forecasting framework, HeatGNN, to improve the forecasting performance. We incorporate the epidemic mechanism model into the epidemiology-informed embedding learning to learn intrinsic epidemic transmission patterns. We use the physics loss to constrain the epidemiology-informed location embeddings to align with real-world epidemics. We propose a mechanistic affinity graph to represent the mechanistic dependence of intrinsic epidemic transmission between locations. We design a heterogeneous transmission graph network to account for the mechanistic heterogeneity. Finally, the forecasting performance is further improved by combining the spatio-temporal graph learning and transmission graph modules. The effectiveness of HeatGNN is validated using real-world datasets for epidemic forecasting. Experimental results demonstrate that HeatGNN is capable of outperforming existing state-of-the-art models in epidemic forecasting. In the future, we plan to construct an interpretable mechanistic affinity graph with mechanisms models to cope with the spatio-temporal and mechanistic heterogeneity.

### v2

@@ VI Conclusion
In this study, we propose a novel epidemic forecasting framework, HeatGNN, to improve the forecasting performance. We incorporate the epidemic mechanism model into the epidemiology-informed embedding learning to learn intrinsic epidemic transmission patterns. We use the physics loss to constrain the epidemiology-informed location embeddings to align with real-world epidemics. We propose a mechanistic affinity graph to represent the mechanistic dependence of intrinsic epidemic transmission between locations. We design a heterogeneous transmission graph network to account for the mechanistic heterogeneity. Finally, the forecasting performance is further improved by combining the spatio-temporal graph learning and transmission graph modules. The effectiveness of HeatGNN is validated using real-world datasets for epidemic forecasting. Experimental results demonstrate that HeatGNN is capable of outperforming existing state-of-the-art models in epidemic forecasting. In the future, we plan to extend HeatGNN to different epidemics and construct a more interpretable mechanistic affinity graph with mechanisms models to cope further with the spatio-temporal and mechanistic heterogeneity in epidemic transmission.
[Uncaptioned image] Yufan Zheng received the B.S. degree from the Nanfang College of Sun Yat-sen University, Guangdong, China. He was a research assistant at the City University of Hong Kong from December 2022 to July 2024. He is a Ph.D. student at the University of Canberra, Australia, and the ARC Training Centre in Plant Biosecurity. His research interests include data-driven modeling, data mining, and computational epidemiology.
[Uncaptioned image] Wei Jiang received the Master of Data Science degree from the University of Queensland. He is a Ph.D. student in the School of Electrical Engineering and Computer Science at the University of Queensland. His research interests include social media data mining, graph representation learning, and the applications of large language models.
[Uncaptioned image] Tong Chen received the Ph.D. degree in computer science from The University of Queensland, Brisbane, QLD, Australia, in 2020. He is currently a Senior Lecturer with the Data Science Research Group, School of Electrical Engineering and Computer Science, The University of Queensland. His research interests include data mining, recommender systems, user behavior modeling, and predictive analytics.
[Uncaptioned image] Alexander Zhou is currently a Research Assistant Professor in the Department of Computing at The Hong Kong Polytechnic University. He received his PhD from The Hong Kong University of Science and Technology in 2024. His research interests include graph databases and algorithms as well as social science-driven research on networks.
[Uncaptioned image] Choujun Zhan received the B.S. degree in automatic control engineering from Sun Yat-sen University, Guangzhou, China, in 2007 and the Ph.D. degree in electronic engineering from the City University of Hong Kong in 2012. He worked as a Postdoctoral Fellow at the Hong Kong Polytechnic University after graduation. Since 2016, he has been an Associate Professor at South China Normal University. His research interests include complex networks, time series modeling, epidemic spreading, information diffusion, and machine learning.
[Uncaptioned image] Quoc Viet Hung Nguyen received the Ph.D. degree from EPFL, Switzerland. He is an associate professor at Griffith University, Australia. His research focuses on data integration, data quality, information retrieval, trust management, recommender systems, machine learning, and big data visualisation. He published several papers in top-tier venues such as SIGMOD, VLDB, WSDM, WWW, SIGIR, KDD, AAAI, ICDE, IJCAI, VLDBJ, IEEE TKDE, and ACM TOIS.
[Uncaptioned image] Hongzhi Yin received the Ph.D. degree in computer science from Peking University, in 2014. He works as an ARC Future Fellow and Full Professor at The University of Queensland, Australia. He has made notable contributions to recommendation systems, graph learning, and decentralized and edge intelligence. He has published 300+ papers with an H-index of 84 and received the ARC Future Fellowship 2021 and DECRA Award 2016.